WEAVE(1) WEAVE(1) NAME weave−translateWEBtoTeX SYNOPSIS weave[−x]webfile[.web][changefile[.ch]] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. TheweaveprogramisusedtocreateaTEXfileforviewingaWEBprogram.Ittakesappropriate careoftypographicdetailslikepagelayoutandtheuseofindentation,italics,boldface,etc.,and itsuppliesextensivecross-indexinformationthatitgathersautomatically.Thecommandline argumentsarethesameasfortangleexceptfortheoption:−xsaystoomittheindex,module namelist,andtableofcontentspages.(ACONTENTS.texfilewillstillbewrittenwhentheTEX fileisprocessed,however,unlesssomemacrosinwebmac.texareredefined.) TheoutputTEXfilenameisformedbyadding.textotherootoftheWEBfilename. Thereareseveralmacrosthatprobablyshouldberedefinedbytheprogrammeratthebeginning oftheWEBfile.Itisagoodideatoset\titletothenameoftheprogram.And,tocauseoutput ofonlychangedmodules,onecansay\let\maybe=\iffalse(usuallyasthefirstchangeinthe changefile). FILES $TEXMFMAIN/plain/base/webmac.tex TEXmacrosusedbyweaveoutput. SEEALSO tangle(1)forreferences,authors,andotherinformation. Web2C7.32x 4May1993 1